Let f(x) = xlnx. Approximate f(2) by the Hermite interpolating polynomial using x0 = 1 and x1 = e and compare the error.(e ≈ 2.7...)
Answer #1

f(x) = xlna f(x) = x1 + enx = 14 dna 20=1 , 91=e. To compute: f(2) fee, 2 ln 2. Hermite Interpolation for mula is; pens = È_~) = 1 1+oX EX (7-1){re xo + {1- 2 x L x (1-1)} x (x-1) re. + (x-1) [42.] 1 + (9-e) (24)?. 2 + = 11 - (2-Uca-e12 Cheiatex-e) firstly we obtained the value of f(2) by Hermite Interpolation and then the exact value and by comparing the two we found the approximate error.